KPre-test Problem 19.13 1 of 14 At what separation is the electrostatic force between a +11.1 μC point charge and a +26.5 pC point charge equal in magnitude to 1.50 N? T-
Answer #1

We know for electrostatic force

force {\displaystyle F={\frac {1}{4\pi \varepsilon _{0}}}{\frac {qQ}{r^{2}}}=k_{0}{\frac {qQ}{r^{2}}}\,,}

{\displaystyle k_{0}\approx {\frac {1}{4\pi \varepsilon _{0}}}\approx 8.987\ 551\ 787\times 10^{9}\;\;\mathrm {N\ m^{2}\ C} ^{-2}.}   

K0  = 9 * 109 N m2C-2

given q= 11.1 micro C = 11.1 * 10-6 C

Q=26.7 micro C =26.7* 10-6 C

F=1.50 N

put on formula above

1.50 = (9 * 109 *11.1 * 10-6 *26.7* 10-6) / r2

r2 = (9 * 109 *11.1 * 10-6 *26.7* 10-6) / 1.50

r=1.33 m

Distace between both is 1.33m